The numerator o a fraction is 6 less than its denominator, and the value of the fraction is 3/5
Answer by fcabanski(1391)

The denominator is x since it's unknown. The numerator (top number) is 6 less than x or x-6. So the fraction is (x-6)/x and it is equal to 3/5


(x-6)/x = 3/5 cross multiply and solve for x: 5x-30 = 3x


2x = 30


x = 15. So the fraction is (15-6)/15 = 9/15 which reduces to 3/5.